OpenStudy (anonymous):

Integral sin(2x) (cos(x))^2 dx Substitution rule o by parts ?

myininaya (myininaya):

sin(2x)=2sin(x)cos(x) use this

myininaya (myininaya):

\[2 \int\limits_{}^{}\sin(x)(\cos(x))^3 dx\]

myininaya (myininaya):

use sub u=cos(x)

myininaya (myininaya):

let me know if you need more help

OpenStudy (anonymous):

thx a lot

myininaya (myininaya):

great! you got the rest? :)

OpenStudy (anonymous):

yeah i guess

myininaya (myininaya):

ok well just let me know if you run into troubles

OpenStudy (anonymous):

k

OpenStudy (anonymous):

the solution is -1/2cos^4 (x) +c

myininaya (myininaya):

\[u=\cos(x)\] => \[ du=-\sin(x) dx\] \[-2 \int\limits_{}^{} u^3 du=-2 \cdot \frac{u^4}{4}+C=-\frac{1}{2} u^4+C=-\frac{1}{2}\cos^4(x)+C\]

myininaya (myininaya):

yes thats right! :)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

gor the help

OpenStudy (anonymous):

for+*